OpenStudy (anonymous):

Find the exponential function f(x) = ax whose graph is given. or atleast if i can make the graph for y'all i'll five ya the points

OpenStudy (anonymous):

 but much neater lol

OpenStudy (anonymous):

at x=0, y=1 and at x=2, y=25 so just plug into the exponential equation: \(\large y=a^x \) \(\large 1=a^0 \) , this is true no matter what a is for a>0 \(\large 25=a^2 \rightarrow a=\pm5\rightarrow a=5\) so your exponential equation is \(\large y=5^x \)

OpenStudy (anonymous):

Thank YOU!!

OpenStudy (anonymous):

yw...:)